OpenStudy (anonymous):

the population of a bacteria culture doubles in number every 12 minutes. the ratio of the number of bacteria at the end of 1 hour to the number of bacteria at the beginning of the hour is?

OpenStudy (anonymous):

Let's say the number of bacteria you initially have is "a." The number of bacteria doubles every twelve minutes, which means that it doubles 5 times in an hour. You have \[(2)(2)(2)(2)(2)a=a*2^5=32a\] at the end of the hour. You started out with only a. This means your ratio is 32a/a = 32.